OpenStudy (anonymous):

Find the equation of the ellipse with the following properties. The ellipse with x-intercepts (5, 0) and (-5, 0); y-intercepts (0, 3) and (0, -3).

OpenStudy (anonymous):

would i plug points into this equation

OpenStudy (anonymous):

@thomaster

OpenStudy (anonymous):

@phi

OpenStudy (mathmale):

DD: Please consider graphing this ellipse. You are given four points on the ellipse. Identify which axis is longer (distance from one vertex to the opposite vertex. The longer half vertex (that is, the distance from the center (origin) to the vertex) is " a " and the shorter one is " b ." Why is this necessary? Because the standard equation of an ellipse is \[\frac{ x^2 }{ a^2 }+\frac{ y^2 }{ b^2 }=1\]

OpenStudy (mathmale):

when the ellipse is "horizontal." If it's vertical, then the std. eq'n. is \[\frac{ y^2 }{ a^2 }+\frac{x^2 }{ b^2 }=1\]

OpenStudy (mathmale):

Determine a and b, and then write out the equation of this ellipse.

OpenStudy (anonymous):

ok thank you for the help if you have any other information that would help its more than welcome....unfortunatly i have to go do something i will be back to test my luck....thank you for the help

OpenStudy (mathmale):

My pleasure.